You need to prepare a 30mL solution of a 1:8 syrup solution. You have on hand a 50% syrup solution and a 1:20 soda solution. How many mL of each solution do you need to create this final solution?
Answer #1

Call the amount of 50% solution to be used, x

Then 30 - x is the amount of 1/200 solution used

The set up will then be,

0.5x +0.05(30 -x) = 1/8 (30)

0.5x + 1.5 -0.05x = 3.75

0.45x = 2.25

X = 5 ml of 50% syrup solution will be used